find the slope (-1,-2) (3,-25) thank you
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

(-1,-2) (3,-25)
Slope=(y2-y1)/(x2-x1)
Slope=(-25+2)/(3-1)
Slope=-23/2 ans.

Answer by Alan3354(69443)

m = diffy/diffx = -23/4
m = -23/4